OpenStudy (anonymous):

help please:) If the sin 75° is 0.966, then the cos _____° is 0.966.

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

help?!!

OpenStudy (jdoe0001):

\(\bf \textit{Cofunction Identities} \\ \quad \\ sin\left(90^o-{\color{red}{ \theta}}\right)=cos({\color{red}{ \theta}})\qquad cos\left(90^o-{\color{red}{ \theta}}\right)=sin({\color{red}{ \theta}})\) what do you think?

OpenStudy (anonymous):

hhhmm,...sorry,...i don't know

OpenStudy (jdoe0001):

lemme put it this way then \(\bf cos\left(90^o-{\color{red}{ \theta}}\right)=sin({\color{red}{ 75^o}})\)

OpenStudy (anonymous):

i am really sorry i still don't get it:(

OpenStudy (jdoe0001):

what about \(\bf cos\left(90^o-{\color{red}{ 75^o}}\right)=sin({\color{red}{ 75^o}}) \) makes any sense?

OpenStudy (triciaal):